OpenStudy (anonymous):

y+x = 5 5x + 5y = 25 Infinite number None One

jhonyy9 (jhonyy9):

are the same functions

OpenStudy (anonymous):

factorize the second equation 5x+5y=25 5(x+y)=25 x+y=25/5 x+y=5 they are the same !

OpenStudy (anonymous):

that means that, for any value x has in this system, there is a y that exists which respects both equations. For example x =2, y =3. x = 4, y = 1. X= 9932, y = -9927

OpenStudy (anonymous):

So the answer is None

OpenStudy (anonymous):

no, the answer is infinite because x can be anything, and y just needs to be 5 - x

OpenStudy (anonymous):

oh!! o.k. I get it now